OpenStudy (anonymous):

how do you factor this: 2y²-9y+4

OpenStudy (anonymous):

(2y-1)(y-4) 2y^2-8y-1y=4 = 2y^2-9y=4

OpenStudy (anonymous):

\[(2y-1)(y-4)=2y ^{2}-8y-1y+4 \]